OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that has a slope of 4 and contains the point (-1,-3)

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\] m=slope (x1,y1)

pooja195 (pooja195):

We know that the slope is 4 so m= slope We know x1 = -1 we know y1=-3 now just plug those into the formula :)

OpenStudy (anonymous):

2

pooja195 (pooja195):

huh? 2 for what?

pooja195 (pooja195):

i plugged it in for you :) now solve \[ \huge~y+3=4(x+1) \]

pooja195 (pooja195):

Do you know what to do after this?

OpenStudy (anonymous):

multiply 1 from 4

OpenStudy (anonymous):

?

pooja195 (pooja195):

no distribute

OpenStudy (anonymous):

how do i distribute

pooja195 (pooja195):

4(x+1) 4*x=? 4*1=?

OpenStudy (anonymous):

4

pooja195 (pooja195):

y+3=4x+4 now we need to isolate y so we need to subtract 4-3=?

OpenStudy (anonymous):

1

pooja195 (pooja195):

good so our final answer would be??

OpenStudy (anonymous):

y=4x+1

pooja195 (pooja195):

good